OpenStudy (chrisneedshelp):

.Solve the system algebraically. 4x - 5y = 3 x = 3 - y What is the value of x?

OpenStudy (shamil98):

Use substitution you are given the value of x. 4x-5y= 3 x = 3-y 4(3-y) - 5y = 3 and solve for y.